OpenStudy (anonymous):

5m of soil containing 50% silt was mixed into 3 m of soil containing 10% silt. What is the silt content of the mixture?

OpenStudy (amistre64):

5(.5) + 3(.1) = 8(x) 2.5 + .3 = 8x 2.8 = 8x 2.8/8 = x 1.4/4 = .7/2 = .35 35%

OpenStudy (anonymous):

thank u